OpenStudy (anonymous):

Total the Riemann sum, for f(x) = 4x + 3 on [1, 4]; n =6.

jimthompson5910 (jim_thompson5910):

You're going to be evaluating this function at values that span (4-1)/6 = 3/6 = 1/2 Basically, you'll start out plugging in x=1 to find f(1) Then you'll add 1/2 to 1 to get x=3/2 and evaluate the same function at this x value to get f(3/2) Then you add another 1/2 to get x=2 and evaluate to find f(2) etc... Keep doing this until you reach x = 4. You'll find that there will be 6 values you'll need to evaluate. Once you have your corresponding function outputs, simply add them up and that sum will be your answer.